what is the standard form equation for points ( 3,5 ) slope equals 5/3

Answer by stanbon(75887)

what is the standard form equation for points ( 3,5 ) slope equals 5/3
------
Form: y = mx + b
5 = (5/3)3 + b
5 = 5 + b
b = 0
----
y = (5/3)x
----
Standard form is ax + by = c
5x - 3y = 0
================================
Cheers,
Stan H.